OpenStudy (anonymous):

There are 5 red and 9 green balls in a bag.

OpenStudy (anonymous):

is that the whole problem

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

5 + 9 = 14 5/14 = 0.35 9/14 = 0.65 35% red 65% green

OpenStudy (anonymous):

how did you know what the question was

OpenStudy (anonymous):

I just assumed he was looking for the ratio of red to green.

OpenStudy (anonymous):

oh okay smart boy